# A NICER view of the â edge-on pulsar PSR J06143329

### v1

@@ 5.1 X-PSI resolution
The effects of the LR settings on the likelihood were investigated by drawing samples from the posterior of the models presented in Section 4 and computing a likelihood with high resolution (HR) settings for each sample, which was then compared against the likelihood with LR (exploratory) or XR (headline) settings used during sampling. The distributions are plotted in Figure 6. The distributions are all compatible with at the level. The spread, which measures the accuracy of the likelihood, increases with model complexity, reaching for PDT-U. This is because, under the same X-PSI resolution settings, we observe a higher likelihood accuracy for ST regions compared to PDT regions, as the former exhibit a simpler geometry. While the accuracy (i.e., an maximum value of ) needed for precise likelihood computation is difficult to determine, we deem these values low compared to the range of values taken by the likelihood during sampling, . Moreover, a log-evidence of would be needed to claim a significant superiority of a model over another, which is times more than our reported accuracy. Hence we could not find any systematic bias between LR and HR likelihoods for the posterior geometries, although we can conclude that PDT-Uâ likelihoods were less accurate than those of the other models.
While more likely, more complex, thin, elongated structures with biased LR likelihoods could be missed while sampling, these distributions of LR runs support that the convergence of their posteriors is consistent with a local maximum of the HR model. It is however not possible from such consideration to conclude that a complex structure was not missed. Hence, the Headline ST+PDTâ model was run with a higher X-PSI resolution and sampling accuracy to account for this possibility and converged toward the same solution as the exploratory run (see Section 4.4).
@@ 5.2 Noise realization
The Poisson noise in the NICER data can have an effect on the inferred parameters (Vinciguerra et al., 2023). This phenomenon has also been investigated here by testing different binning of the phases for the PSR J06143329 data. Even though the phase measurement of all collected photons is very precise (s), the data used for likelihood computation (Figure 1) is binned in rotation cycles (about s wide) bins to save on computational time. By uniformly shifting the phase of individual NICER photons by a small amount ( bin) before binning, most photons will remain in the same bins, and some will end up in the neighboring bin, therefore resulting in slightly different bolometric pulse profiles. This way, we mimic different noise realizations of the binned pulse profile, which however all share the same spectrum and XMM data. Using shifts of and cycles, that are smaller than the bin width, we ran LR runs with the preferred ST+PDTâ model.
@@
The log-evidences and maximum likelihoods have strong dependence on the binning of the data. This dependence (given in Table 1) is greater than the difference in metrics between all the ST-U, ST+PDTâ and PDT-Uâ models.
@@ 5.3 Model comparison
As it was not trivial to differentiate ST+PDTâ and PDT-Uâ from log-evidence values alone (see Section 4.3), we turned to the arviz package (Vehtari et al., 2015) which provides model comparison techniques taking into account the complexity of the models and comparing their predictive power. This requires a point-wise likelihood, i.e., a likelihood value for each individual bin. This product is however not available in our case, since our inferences use channel-wise background marginalized likelihoods, which is not applicable bin by bin (see Section 3.5). Instead, we use a simple Poisson likelihood and specify as supplementary input parameters the background parameter values.
@@ 6.1 Comparison to other MSPs with PPM constraints
The inferred radius of PSR J06143329 is constrained at the level ( credible interval)666These numbers do not include systematic uncertainties, which are difficult to quantify. A rough estimate can be obtained by looking at the dispersion of radius values for different models. For instance, it is of the order of for PSR J06143329 and for PSR J00300451., which is not as tight as for PSR J04374715 (about , Choudhury et al. 2024a), but is of the order of PSR J07406620 (about , Salmi et al. 2024a). The radius constraint of PSR J12311411 (Salmi et al., 2024b) is harder to compare, since that study used different radius priors, some of which were EOS informed, causing the tighter constraint. Some individual models for PSR J00300451 achieved lower uncertainty thanks to a higher photon statistics, resulting from a deeper exposure and a brighter source (Riley et al., 2019; Vinciguerra et al., 2024).
While Vinciguerra et al. (2024) incorporated background constraints in the PSR J00300451 analysis, Riley et al. (2019) did not. The former resulted in different models fitting the data equally well, partially because of the lack of a mass prior. Our result is consistent with the ST+PDTâ solution for PSR J00300451 from Vinciguerra et al. (2024). However, our inferred radius is not directly comparable to that of the PDT-Uâ solution from PSR J00300451, as this solution points toward a mass of 1.7â instead of 1.4â.
Interestingly, the measured radius is compatible at the level with the inferred radius of PSR J04374715, a MSP of a similar mass (Choudhury et al., 2024a) and with the indirectly inferred777GW events carry information on tidal deformability. Extracting radius information from this information is an EOS-dependent process. mass-radius contours for the NSs whose merger produced the GW170817 signal (Abbott et al., 2018).
@@
PSR J04374715 is very bright with a tight prior on its mass, inclination and distance (Reardon et al., 2024) and also has little change in measured radius across models (Choudhury et al., 2024a), but an AGN is in the field of view and a 3C50-informed background was used, both introducing possible bias. Moreover, PSR J04374715 is seen at a inclination, which may allow for a completely different hotspot configuration when changing models. This does not seem to be the case for PSR J06143329, which is seen edge-on. Overall, PSR J06143329 likely owes its radius measurement stability to its tight prior from radio timing, along with the fact that it is seen edge-on.
@@ 6.2 Implications of PSRÂ J06143329 on the dense matter EOS
With the new PPM-derived mass-radius posteriors of PSR J06143329 in hand, a subsequent EOS inference is necessary to understand the impacts of this new measurement on the inferred dense matter EOS and NS properties. To perform an EOS inference, we build on the work of Rutherford et al. (2024). Starting from the Baym-Pethick-Sutherland crust EOS (BPS, Baym et al. 1971) for densities below , where is the nuclear saturation density, we use the EOS band from next-to-next-to-next-to-leading order (N3LO) chiral effective field theory (EFT) calculations up to (Keller et al., 2023). Beyond this density, we use two different high-density EOS parameterizations: one based on a piecewise-polytropic model (PP, Hebeler et al. 2013) and one based on the speed of sound in the NS (CS, Greif et al. 2019). This results in a very large prior range for the EOS. For details on how the dense matter EOS models are constructed and their prior assumptions, we refer to Rutherford et al. (2024, and references therein).
@@
We obtain posterior EOS distributions through sampling with MultiNest. This entails first evaluating the priors based on allowed EOSs in the EFT range and for a broad range of parameters that span the PP or CS extensions. The likelihood incorporates the mass-radius constraints of MSPs obtained with PPM and mass-tidal deformability from GW events. This framework is implemented using the open-source dense-matter EOS inference code NEOST888https://github.com/xpsi-group/neost (v2.1.0, Raaijmakers et al. 2025). NEOST has been used for NICER EOS constraints and other dense matter inferences with prerelease versions having been employed in Greif et al. (2019); Raaijmakers et al. (2019, 2020, 2021); Rutherford et al. (2023).
@@
We compare our new EOS inferences to the âNewâ scenario from Rutherford et al. (2024). This scenario is defined by the combination of the ST-U solution of PSR J07406620 from Salmi et al. (2024a), the ST+PDT solution of PSR J00300451 from Vinciguerra et al. (2024), the CST+PDT solution of PSR J04374715 from Choudhury et al. (2024a), and the mass-tidal deformability posteriors of GW170817 and GW190425 from Abbott et al. (2019) and Abbott et al. (2020a), respectively. We investigate the effect of adding the new mass-radius result of PSR J06143329 to this data set. Interestingly, with the new result from PSR J06143329 included, the EOS inference results would be very similar if only MSP with mass and inclination priors were accounted for and PSR J00300451 (ST+PDT) left out (see the Zenodo repository for figures and samples of the inferred EOS in that case).
@@
The EOS inference yields mass-radius posterior distributions (Figure 9) centered around 12âkm with maximum masses for both the PP and CS extensions. Our results for the radii of 1.4 and NSs, as well as the maximum mass configurations are also reported in Table 2. Compared to Rutherford et al. (2024) (their âNewâ scenario), the mass-radius posteriors shift toward lower radii by m, with a minimal impact on the maximum mass for both PP and CS models. This slight shift is not only due to the addition of PSR J06143329, but also to the fact that the inferred radius of PSR J06143329 is consistent with those of PSR J00300451 (ST+PDT) and PSR J04374715.
@@
At the 68% credible interval, the PP model posteriors show a bimodal-like structure. This is also seen in both the PP and CS models for the distribution of the radius of a NS in the insets of Figure 9. This bimodal-like structure is already present in the posterior mass-radius distributions without the inclusion of PSR J06143329, but is less pronounced for the PP model and evenly peaked below and above 12âkm for the CS model (see Figure 7 of Rutherford et al., 2024). These differences in the bimodal-like structure are the result of including PSR J06143329, which offers more posterior support for mass-radius samples below 12âkm, making this slightly less pronounced for the CS model and more pronounced for the PP model. This suggests that the inclusion of PSR J06143329 slightly softens the EOS. This trend is confirmed in the pressure posterior distributions at and shown in Figure 10 (see also the values in Table 2), where the pressure is pushed to slightly lower values, while still being on the stiffer end of the priors, confirming a pattern seen in Raaijmakers et al. (2021) and Rutherford et al. (2024). As expected, the bimodal-like feature is also present in the pressure posterior distributions, particularly at .
@@
The EOS slightly softening with the inclusion of PSR J06143329 is further confirmed by the difference of the radii of 2.0 and NSs, . Adding PSR J06143329 allows for slightly more negative values, which is expected, given that the inclusion of PSR J06143329 pushes the mass-radius posteriors to smaller radii, particularly for NSs around . These values, along with other key quantities, are provided on Table 2.
@@ 6.3 Caveats
During our analyses, we did not try the Concentric Single Temperature (CST) or Concentric Double Temperature (CDT) hot region models (see Riley et al. 2019). These are made of 2 concentric overlapping circular spots of different temperatures. The configuration of such models can be reproduced by the PDT model, where both spot centers align, which would result in a parameter (angular distance between ceding and superseding components centers) to be close to . For the PDT-Uâ model, we found for both hot regions, while for the Headline ST+PDT, the mode closest to such a configuration had . In the latter case, although is small it is only compatible with 0 at .
@@
Although the PST, CST and CDT geometric models did not appear as specific solutions of a PDT region, it would however be interesting to test them in further analyses of this source.
This issue of model choice is central in PPM, as it affects inferred radii constraints (see for instance Vinciguerra et al. 2024). Choosing an overly complex model can result in biased radius measurements with overestimated evidence, as it might be the case with PDT-U. Here, we chose for our headline result to use the ST+PDTâ model over the PDT-Uâ one which has better evidence, although not significantly. Thorough model comparison was undertaken, but we could not definitively tell whether the PDT-Uâ model was preferred or not (see Section 5.3). The similarity and consistency of radii and geometry constraints from PDT-Uâ and ST+PDT, as well as ST+PDTâ being the simplest model fitting the data well, motivated the choice of the latter for our headline run. Our limited computation resources also weighed in on this choice. Further X-PSI analyses of this source could be carried out with a high resolution PDT-Uâ model.
Although many tests have been done in Section 5 to assess the efficiency of the sampling and the accuracy of the likelihood, it is impossible to assert the completeness of the sampling. There is, however, no strong evidence supporting the incomplete sampling case and the procedure seems to have effectively converged as the headline ST+PDTâ run yielded similar results as the exploratory run. Moreover, we also showed that the LR exploratory run posterior samples have likelihoods consistent with that of a high resolution likelihood (Section 5.1) and that our inference is resilient to noise (Section 5.2).

### v2

@@ 5.1 X-PSI resolution
The effects of the LR settings on the likelihood were investigated by drawing samples from the posterior of the models presented in Section 4 and computing a likelihood with high resolution (HR) settings for each sample, which was then compared against the likelihood with LR (exploratory) or XR (headline) settings used during sampling. The distributions are plotted in Figure 6. The distributions are all compatible with at the level. The spread, which measures the accuracy of the likelihood, increases with model complexity, reaching for PDT-U. This is because, under the same X-PSI resolution settings, we observe a higher likelihood accuracy for ST regions compared to PDT regions, as the former exhibit a simpler geometry. While the accuracy (i.e., a maximum value of ) needed for precise likelihood computation is difficult to determine, we deem these values low compared to the range of values taken by the likelihood during sampling, . Moreover, a log-evidence difference of would be needed to claim a significant superiority of a model over another, which is times more than our reported accuracy. Hence we could not find any systematic bias between LR and HR likelihoods for the posterior geometries, although we can conclude that PDT-Uâ likelihoods were less accurate than those of the other models.
Even if their likelihoods under HR settings are higher, configurations with more complex, thin, elongated structures with biased LR settings likelihoods could be missed while sampling. However, these distributions of LR runs support LR being accurate in the posterior space of geometries for all ST-U, ST+PDTâ and PDT-U, and their posteriors are consistent with a local maximum of the HR model. Nevertheless, we did not demonstrate that LR settings yield accurate likelihoods over the whole prior space of geometries and it is hence not possible from such considerations to conclude that a complex structure was not missed. It is possible that, somewhere in the prior space, there is a better fitting geometric configuration for which LR likelihoods are not accurate and biased toward lower values. Such a region could be unknowingly missed during sampling and would appear when using high X-PSI resolution. The Headline ST+PDTâ model was run with a higher X-PSI resolution and sampling accuracy to account for this possibility and converged toward the same solution as the exploratory run (see Section 4.4). This indicates that it could also be the case for PDT-U, which can be confirmed with a higher resolution PDT-Uâ run.
@@ 5.2 Noise realization
The Poisson noise in the NICER data can have an effect on the inferred parameters (vinciguerra_x-psi_2023). This phenomenon has also been investigated here by testing different binning of the phases for the PSR J06143329 data. Even though the phase measurement of all collected photons is very precise (s), the data used for likelihood computation (Figure 1) is binned in rotation cycles (about s wide) bins to save on computational time. By uniformly shifting the phase of individual NICER photons by a small amount ( bin) before binning, most photons will remain in the same bins, and some will end up in the neighboring bin, therefore resulting in slightly different bolometric pulse profiles. This way, we mimic different noise realizations of the binned pulse profile, which, however, all share the same spectrum and XMM data. Using shifts of and cycles, that are smaller than the bin width, we ran LR runs with the preferred ST+PDTâ model.
@@
The log-evidences and maximum likelihoods have strong dependence on the binning of the data. This dependence (given in Table 1) is greater than the difference in metrics between all the ST-U, ST+PDT, and PDT-Uâ models.
@@ 5.3 Model comparison
As it was not trivial to differentiate ST+PDTâ and PDT-Uâ from log-evidence values alone (see Section 4.3), we turned to the arviz package (vehtari_practical_2017), which provides model comparison techniques taking into account the complexity of the models and comparing their predictive power. This requires a point-wise likelihood, i.e., a likelihood value for each individual bin. This product is, however, not available in our case, since our inferences use channel-wise background marginalized likelihoods, which is not applicable bin by bin (see Section 3.5). Instead, we use a simple Poisson likelihood and specify as supplementary input parameters the background parameter values.
@@ 6.1 Comparison to other MSPs with PPM constraints
The inferred radius of PSRÂ J06143329 is constrained at the level ( credible interval)555These numbers do not include systematic uncertainties, which are difficult to quantify. A rough estimate can be obtained by looking at the dispersion of radius values for different models. For instance, it is of the order of for PSRÂ J06143329 and for PSRÂ J00300451., which is not as tight as for PSRÂ J04374715 (about , choudhury_nicer_2024), but is of the order of PSRÂ J07406620 (about , salmi_radius_2024). The radius constraint of PSRÂ J12311411 (salmi_J1231_2024) is harder to compare, since that study used different radius priors, some of which were EOS informed, causing the tighter constraint. Some individual models for PSRÂ J00300451 achieved lower uncertainty thanks to higher photon statistics, resulting from a deeper exposure and a brighter source (riley_nicer_2019; vinciguerra_updated_2024).
While vinciguerra_updated_2024 incorporated background constraints in the PSRÂ J00300451 analysis, riley_nicer_2019 did not. The former resulted in different models fitting the data equally well, partially because of the lack of a mass prior. Our result is consistent with the ST+PDTâ solution for PSRÂ J00300451 from vinciguerra_updated_2024. However, our inferred radius is not directly comparable to that of the PDT-Uâ solution from PSRÂ J00300451, as this solution points toward a mass of 1.7â instead of 1.4â.
Interestingly, the measured radius is compatible at the level with the inferred radius of PSRÂ J04374715, a MSP of a similar mass (choudhury_nicer_2024) and with the indirectly inferred666GW events carry information on tidal deformability. Extracting radius information from this information is an EOS-dependent process. mass-radius contours for the NSs whose merger produced the GW170817 signal (abbott_gw170817_2018).
@@
PSRÂ J04374715 is very bright with a tight prior on its mass, inclination and distance (reardon_neutron_2024) and also has little change in measured radius across models (choudhury_nicer_2024), but an AGN is in the field of view and a 3C50-informed background was used, both introducing possible bias. Moreover, PSRÂ J04374715 is seen at a inclination, which may allow for a completely different hotspot configuration when changing models. This does not seem to be the case for PSRÂ J06143329, which is seen edge-on. Overall, PSRÂ J06143329 likely owes its radius measurement stability to its tight prior from radio timing, along with the fact that it is seen edge-on.
@@ 6.2 Implications of PSRÂ J06143329 on the dense matter EOS
With the new PPM-derived mass-radius posteriors of PSRÂ J06143329 in hand, a subsequent EOS inference is necessary to understand the impacts of this new measurement on the inferred dense matter EOS and NS properties. To perform an EOS inference, we build on the work of rutherford_constraining_2024. Starting from the Baym-Pethick-Sutherland crust EOS (Baym71) for densities below , where is the nuclear saturation density, we use the EOS band from next-to-next-to-next-to-leading order (N3LO) chiral effective field theory (EFT) calculations up to (Keller2023). Beyond this density, we use two different high-density EOS parameterizations: one based on a piecewise-polytropic model (PP, Hebeler2013) and one based on the speed of sound in the NS (CS, Greif19). This results in a very large prior range for the EOS. For details on how the dense matter EOS models are constructed and their prior assumptions, we refer to rutherford_constraining_2024.
@@
We obtain posterior EOS distributions through sampling with MultiNest. This entails first evaluating the priors based on allowed EOSs in the EFT range and for a broad range of parameters that span the PP or CS extensions. The likelihood incorporates the mass-radius constraints of MSPs obtained with PPM and mass-tidal deformability from GW events. This framework is implemented using the open-source dense matter EOS inference code NEOST777https://github.com/xpsi-group/neost (v2.1.0, Raaijmakers24). NEOST has been used for NICER EOS constraints and other dense matter inferences with prerelease versions having been employed in Greif19; Raaijmakers19; Raaijmakers20; Raaijmakers21; Rutherford:2022xeb.
@@
We compare our new EOS inferences to the âNewâ scenario from rutherford_constraining_2024. This scenario is defined by the combination of the ST-U solution of PSRÂ J07406620 from salmi_radius_2024, the ST+PDT solution of PSRÂ J00300451 from vinciguerra_updated_2024, the CST+PDT solution of PSRÂ J04374715 from choudhury_nicer_2024, and the mass-tidal deformability posteriors of GW170817 and GW190425 from PhysRevX.9.011001 and abbott_gw190425_2020, respectively. We investigate the effect of adding the new mass-radius result of PSRÂ J06143329 to this data set. Interestingly, with the new result from PSRÂ J06143329 included, the EOS inference results would be very similar if only MSPs with mass and inclination priors were accounted for and PSRÂ J00300451 (ST+PDT) left out (see the Zenodo repository of zenodo for figures and samples of the inferred EOS in that case).
@@
The EOS inference yields mass-radius posterior distributions (Figure 9) centered around 12âkm with maximum masses for both the PP and CS extensions. Our results for the radii of 1.4 and NSs, as well as the maximum mass configurations are also reported in Table 2. These inferred mass-radius posteriors are less than one standard deviation away from rutherford_constraining_2024 âNewâ data scenario, although they shift toward lower radii by m, with a minimal impact on the maximum mass for both PP and CS models. This slight shift is not only due to the addition of PSR J06143329, but also to the fact that the inferred radius of PSR J06143329 is consistent with those of PSR J00300451 (ST+PDT) and PSR J04374715.
@@
At the 68% credible interval, the PP model posteriors show a bimodal-like structure. This is also seen in both the PP and CS models for the distribution of the radius of a NS in the insets of Figure 9. This bimodal-like structure is already present in the posterior mass-radius distributions without the inclusion of PSR J06143329, but is less pronounced for the PP model and evenly peaked below and above 12âkm for the CS model (see Figure 7 of rutherford_constraining_2024). These bimodal-like distributions are the result of larger radii (stiffer) EOS being favored by both the prior and the large measured mass of PSR J07406620 plus smaller radii (softer) EOS being favored by both PSR J04374715, PSR J06143329, GW170817 data, and, marginally, the ST+PDT contour of PSR J00300451, which overlaps with the mass-radius contours of the latter sources. Including PSR J06143329, which offers more posterior support for mass-radius samples below 12âkm, make this bimodal-like structure slightly less pronounced for the CS model and more pronounced for the PP model. This suggests that the inclusion of PSR J06143329 slightly softens the EOS. This trend is confirmed in the pressure posterior distributions at and shown in Figure 10 (see also the values in Table 2), where the pressure is pushed to slightly lower values, while still being on the stiffer end of the priors, confirming a pattern seen in Raaijmakers21 and rutherford_constraining_2024. As expected, the bimodal-like feature is also present in the pressure posterior distributions, particularly at .
@@
Adding PSR J06143329 allows for slightly more negative difference of the radii of 2.0 and NSs, . This can be explained by a softening of the EOS, as it would reduce more prominently the posterior radius around 2.0  than at 1.4 . This effect can be seen on the lower radius branch of the bimodal-like structure for PP. Hence, the EOS slightly softening with the inclusion of PSR J06143329 is further confirmed. These values, along with other key quantities, are provided on Table 2.
@@
âą
Notes - and are the radii of a ââ and ââNS, with . and are the mass and radius of a non-rotating, maximum mass NS, with , and being its central energy density, density, and pressure, respectively. We also give the pressure at 2, 3, and (see Figure 10 for associated distributions). For this table only, to compare with rutherford_constraining_2024, median values with equal-tailed credible intervals are given.
In addition, if instead of the Headline ST+PDT, the PDT-Uâ geometry had been used for PSR J06143329, we expect that the new EOS inferences would still have stronger support for softer EOS than the âNewâ scenario in rutherford_constraining_2024, while the reported mass-radius contours at and confidence intervals would change notably less from the âNewâ scenario than in the ST+PDTâ case, as mass-radius posteriors from the exploratory PDT-Uâ run mostly overlap the PSR J04374715 ones (Figure 3).
@@ 6.3 Caveats
During our analyses, we did not try the Concentric Single Temperature (CST) or Concentric Double Temperature (CDT) hot region models (see riley_nicer_2019). These are made of two concentric overlapping circular spots of different temperatures. The configuration of such models can be reproduced by the PDT model, where both spot centers align, which would result in a parameter (angular distance between ceding and superseding components centers) to be close to . For the PDT-Uâ model, we found for both hot regions, while for the Headline ST+PDT, the mode closest to such a configuration had . In the latter case, although is small it is only compatible with 0 at .
@@
Although the PST, CST and CDT geometric models did not appear as specific solutions of a PDT region, it would, however, be interesting to test them in further analyses of this source.
This issue of model choice is central in PPM, as it affects inferred radii constraints (see for instance vinciguerra_updated_2024). Choosing an overly complex model can result in biased radius measurements with overestimated evidence, as it might be the case with PDT-U. Here, we chose for our headline result to use the ST+PDTâ model over the PDT-Uâ one, which has better evidence, although not significantly. Thorough model comparison was undertaken, but we could not definitively tell whether the PDT-Uâ model was preferred or not (see Section 5.3). The similarity and consistency of radii and geometry constraints from PDT-Uâ and ST+PDT, as well as ST+PDTâ being the simplest model fitting the data well, motivated the choice of the latter for our headline run. Our limited computation resources also weighed in on this choice. Further X-PSI analyses of this source could be carried out with a high resolution PDT-Uâ model.
Although many tests have been done in Section 5 to assess the efficiency of the sampling and the accuracy of the likelihood, it is impossible to assert the completeness of the sampling. There is, however, no strong evidence supporting the incomplete sampling case and the procedure seems to have effectively converged as the headline ST+PDTâ run yielded similar results as the exploratory run. In retrospect, this supports our choice of LR for exploratory runs (Section 3.1), which proved to be a sufficiently high resolution as we found no discrepancies between the exploratory and Headline ST+PDTâ runs. Using higher resolution X-PSI settings for the Headline ST+PDTâ was, however, required to look for complex geometries that could have been missed during sampling. These higher resolution X-PSI settings were picked judiciously so that computation time was not too heavily affected while improving accuracy by one order of magnitude (Appendix A). Moreover, we also showed that the LR exploratory run posterior samples have likelihoods consistent with that of a high resolution likelihood (Section 5.1) and that our inference is resilient to noise (Section 5.2).